OpenStudy (anonymous):

use the method of substitution to solve the system x^2+y^2=20 x+y=1

OpenStudy (anonymous):

\[x ^{2} + y^{2}=20,\] \[x=1-y\] \[(1-y)^{2}+y^{2}=20\] \[ 1-2y -y^{2} +y^{2}=20 \] \[1-2y=20\] \[-2y=19\] \[y=-19/2\]

OpenStudy (anonymous):

x=1+19/2

OpenStudy (anonymous):

Careful: (1 - y)(1 -y) = 1 - 2y + y^2

OpenStudy (anonymous):

Yeah, i know that is equals 1-2y=y^2 and i tried to factor and even complete the square but I am still baffled